In solving the rational equation x/x^2-7x+12-3/x^2-9=1 the largest non-permissible value of x is?
Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52915)

.
Find the roots of the denominators polynomials and select the maximal value of them.


Answer by stanbon(75887)

In solving the rational equation x/x^2-7x+12-3/x^2-9=1 the largest non-permissible value of x is?
-------
x/(x^2-7x+12) -3/(x^2-9) = 1
-----
Factor the denominators::
x/[(x-4)(x-3)] - 3/[(x+3)(x-3)] = 1
--------------------------------------
Ans: x = 4
-----------------
Cheers,
Stan H.
----------------